The H+2 ion is composed of two protons, each of charge +e=1.60 x 10^-19 C, and an electron of charge -e And mass 9.11 x10^-31 kg. The separation between the protons is 1.07x10^-10 m. The protons and the electron may be treated as point charges.

Suppose the electron in part A has a velocity of magnitude 2.30
Answer #1

a) the potential energy between two points is given by PE = k*Q1*Q2/d

so,

the potential of the electron and protons is = 2*9*10^9*(1.6*10^-19)*(-1.6*10^-19)/(0.5*1.07*10^10^-10)

= - 8.61*10^-18 J

b) We need to conserve energy to get the results.

Initial potential energy + initial kinetic energy = final potential energy

or -8.61*10^-18 + 0.5*9.1*10^-31*(2.3*10^6)^2 = -2*9*10^9*(1.6*10^-19)^2/d

or d=7.43*10^-11 m

so distance from the mid-point be x.so,

x^2 + (0.5*1.07*10^-10)^2=(7.43*10^-11)^2

or x=5.155*10^-11 m